# SynTable: A Synthetic Data Generation Pipeline for Unseen Object Amodal Instance Segmentation of Cluttered Tabletop Scenes

### v2

@@ Abstract
In this work, we present SynTable, a unified and flexible Python-based dataset generator built using NVIDIA’s Isaac Sim Replicator Composer for generating high-quality synthetic datasets for unseen object amodal instance segmentation of cluttered tabletop scenes. Our dataset generation tool can render a complex 3D scene containing object meshes, materials, textures, lighting, and backgrounds. Metadata, such as modal and amodal instance segmentation masks, occlusion masks, depth maps, bounding boxes, and material properties can be generated to annotate the scene according to the users’ requirements automatically. Our tool eliminates the need for manual labeling in the dataset generation process while ensuring the quality and accuracy of the dataset. In this work, we discuss our design goals, framework architecture, and the performance of our tool. We demonstrate the use of a sample dataset generated using SynTable by ray tracing for training a state-of-the-art model, UOAIS-Net. The results show significantly improved performance in Sim-to-Real transfer when evaluated on the OSD-Amodal dataset. We offer this tool as an open-source, easy-to-use, photorealistic dataset generator for advancing research in deep learning and synthetic data generation. The source code for our work is available at https://github.com/ngzhili/SynTable. The links to our demonstration video and our sample dataset can be found in the supplementary materials.
@@ 1 Introduction
Figure 1: (a) Sample RGB outputs of photorealistic cluttered tabletop scenes generated by SynTable pipeline. (b) Visualization of RGB Images, Depth Images, Object Amodal Masks, Object Visible Masks, Object Occlusion Masks, and Object Visible Bounding Boxes. This figure is best viewed zoomed in.
Amodal completion is a perceptual ability that enables the perception of whole objects, even when they are partially occluded [16, 1]. Humans are capable of “filling in” the occluded appearance of invisible objects, owing to their vast experience in perceiving countless objects in various contexts and scenes. This ability to perceive an entire object based on its partial appearance is crucial in enabling reliable and accurate planning of subsequent actions. In modern autonomous robotic systems, comprehending objects in an environment is essential for numerous applications. In vision-based robotic grasping, deep learning perception algorithms are important in extracting critical information about objects in a scene, including their category, shape, appropriate grasping points, and occlusion information. In robotic grasping systems, the ability to infer an occluded object’s complete structure from its visible components allows the robot to plan the grasp order and safely grasp novel occluded objects in cluttered scenes with greater precision and robustness.
There are three acute robotic perception challenges related to amodal instance segmentation: Firstly, the lack of large-scale, high-quality datasets for unseen object amodal instance segmentation (UOAIS) prevents some vision-based grasping systems from achieving their best performance [3]. There are several well-known grasp generation datasets [7, 15, 17, 29, 11]. Still, very few datasets focus on UOAIS [2]. This could be because such datasets are challenging to annotate manually and accurately. While it is possible to annotate grasp poses manually [5, 15, 30], human-annotated amodal instance segmentation masks [2] could be highly subjective and prone to errors. These human annotations are therefore considered imperfect ground truths.
Secondly, although it is possible to generate large synthetic datasets in a simulation environment [28], the issue of visual domain mismatch results in a poor Sim-to-Real transfer that will inevitably reduce the performance of algorithms in real-world applications. This could be because of a lack of domain randomization or because the software generates non-photorealistic scenes for synthetic data.
Thirdly, there is currently a lack of tools available to automatically generate amodal instance segmentation mask annotations to train the amodal completion capability of robotic grasping systems. There is also no effective metric to accurately evaluate a perception model’s ability to determine the occlusion order relationship of objects in a scene. The capacity for amodal completion and adequate comprehension of occlusion relationships between objects within a given scene are crucial for robots working in highly cluttered environments. Amodal completion can aid the planning of pick sequences and avoid collisions with other objects in grasping tasks. However, creating amodal annotations is a task that demands significantly more effort and skill from human annotators than creating modal annotations. Consequently, obtaining amodal annotations through manual annotation is an extremely time-consuming and costly process. For this reason, generating amodal annotations using simulation tools is a much more desirable option.
@@
In this work, we aim to address these three challenges by building a unified and flexible Python-based tool to generate customizable, high-quality, and photorealistic datasets for UOAIS of cluttered tabletop scenes. We focus our work on cluttered tabletop scenes as tabletops are a common place where grasping and manipulation tasks are performed. We develop a novel unified framework that combines the rendering and annotation of scenes and objects in a dataset into a single pipeline. Users can customize the type of scene rendered, the number and variety of objects in the scene, and the types of annotations required for their dataset. Our tool is built on NVIDIA’s Isaac Sim Replicator Composer to leverage its high-fidelity graphics rendering capability.
@@
    We develop a pipeline to automatically render photorealistic cluttered tabletop scenes and generate ground truth amodal instance segmentation masks. This eliminates the need for any manual labeling in the dataset generation process while ensuring the quality and accuracy of the dataset. Based on our pipeline, we designed a dataset generation tool to create photorealistic and accurately-labeled custom datasets for UOAIS (refer to Figure 1(a)). Our tool can be easily added to NVIDIA Isaac Sim to leverage its ability to render complex and photorealistic 3D scenes with high fidelity.
@@
    Our tool provides a rich set of annotations related to amodal instance segmentation (refer to Figure 1(b)): modal (visible) and amodal instance segmentation masks, occlusion masks, occlusion rates, and occlusion order adjacency matrix. Users can easily select which annotations to include in their dataset based on the requirements of their application.
@@
    We generate a large-scale sample synthetic dataset using our tool consisting of amodal instance segmentation labels for users to train and evaluate their models on novel objects. Our dataset will be made publicly available.
@@ 2 Related Works
2.1 Amodal Instance Segmentation in Robotics
In recent years, there has been growing interest in developing amodal instance segmentation methods to enable more robust object detection and tracking in complex scenes. However, several limitations have been encountered in research on amodal instance segmentation in robotic grasping, such as the dearth of large-scale, high-quality training data and visual domain mismatch resulting in poor Sim-to-Real transfer.
Lack of Large-scale High-quality Training Data. Recently, a number of amodal datasets [10, 8, 31] have been developed for indoor scenes. However, there is a lack of high-quality datasets designed for amodal instance segmentation related to robotic grasping and manipulation tasks in cluttered tabletop scenes. Gilles et al. [13] recently presented a new benchmark dataset for evaluating the performance of ambidextrous grasping systems in bin-picking scenarios. Nevertheless, their work is limited in terms of scene diversity and object variety and may not be widely applicable beyond the bin-picking scenario. Richtsfeld et al. [23] proposed the Object Segmentation Database (OSD) to facilitate the development of object segmentation algorithms in robotics and computer vision systems. Suchi et al. [24] designed a semi-automatic labeling tool to produce the Object Cluttered Indoor Dataset (OCID). Despite these efforts, neither OSD nor OCID includes amodal instance segmentation masks in their annotations. Recently, Back et al. [2] introduced amodal annotations to OSD. Nevertheless, this was done manually, which could be a time-consuming and error-prone process.
Sim-to-Real Problem. Xie et al. [28] created the Tabletop Object Dataset (TOD), a dataset comprising scenes that lack photorealism. This characteristic, coupled with the lack of domain randomization, might result in a significant Sim-to-Real gap. The UOAIS-Sim Dataset [2] also shares this limitation, as it lacks domain randomization as well.
@@ 2.2 Tools for Generating Synthetic Datasets
With the rapid development of deep learning, the demand of researchers for synthetic datasets has increased in recent years, which has led to the rise of the development of various tools for generating these datasets [27]. For robotics and computer vision applications, PyBullet and MuJoCo [26] are commonly used physical simulators for generating synthetic data. Xie et al. [28] pre-trained an RGB-D unseen object instance segmentation model using PyBullet. Tobin et al. [25] used MuJoCo to generate synthetic images with domain randomization, which can bridge the Sim-to-Real gap by realistically randomizing 3D content. Simulation tools such as PyBullet and MuJoCo typically come with renderers that are accessible and flexible, but they lack physically based light transport simulation, photorealism, material definitions, and camera effects.
@@ 3 Method
Our dataset generation pipeline is illustrated using the diagram in Figure 2. We use a YAML file to store the parameters and configurations of the scenes to be rendered. The objects and settings required to render the scene are retrieved based on the instructions in the configuration file. We collectively term the objects, materials, and light sources used in our pipeline as assets. Thereafter, the tabletop scene with objects floating above the table is rendered in Isaac Sim. We then run a physical simulation to drop the rendered objects onto the table. For every view within a scene, camera viewpoints and lighting conditions are re-sampled. Subsequently, ground truth annotations are obtained and systematically recorded to create the dataset.
@@ 3.3 Sampling of Camera Viewpoints
To capture annotations for each scene from multiple viewpoints, we enhance the approach by Gilles et al. [18]—which only uses fixed viewpoint positions—by introducing a feature that captures $V$ number of viewpoints at random positions within two concentric hemispheres, as illustrated in Figure 4. $V$ can be set by the user. The radii of the two concentric hemispheres are uniformly sampled within the range $r_{view\_lower}$ m to $r_{view\_upper}$ m, where $r_{view\_lower}$ and $r_{view\_upper}$ are defined in Equations 1 and 2. Users may also set fixed values for $r_{view\_lower}$ and $r_{view\_upper}$ should they wish to do so.
$r_{view\_lower}=max\left(\frac{w}{2},\frac{l}{2}\right)$ (1)
$r_{view\_upper}=1.7\times r_{view\_lower}$ (2)
The hemisphere’s spherical coordinates are parameterized using three variables $r_{view}$, $u$, and $v$. To generate the camera coordinates in the world frame, we first obtain the radius of the hemisphere $r_{view}$ by uniform sampling between $r_{view\_lower}$ and $r_{view\_upper}$. Next, we uniformly sample $u,v$in [0,1], then substitute all the sampled values into Equations 3, 4 and 5 to compute the cartesian coordinates of the camera.
$x=r_{view}\sin(\arccos(1-v))\cos(2\pi u)$ (3)
$y=r_{view}\sin(\arccos(1-v))\sin(2\pi u)$ (4)
$z=r_{view}\cos(\arccos(1-v))$ (5)
@@ 3.4 Sampling of Lighting Conditions
$r_{light\_lower}=r_{view\_upper}+0.1m$ (6)
$r_{light\_upper}=r_{light\_lower}+1m\vspace{-2mm}$ (7)
Figure 6: The process of capturing the ground truth annotations for a scene. For each viewpoint, we capture the following: (a) RGB and depth with all objects displayed, (b) object visible masks & bounding box, (c) object amodal masks, (d) object occlusion masks and occlusion rate, (e) occlusion order adjacency matrix. This figure is best viewed zoomed in.
@@ 3.5 Capturing of Ground Truth Annotations
The process of capturing the ground truth annotations for a scene is illustrated in Figure 6. At each view, the RGB and depth images of the tabletop scene will be captured (Figure 6(a)). The built-in instance segmentation function in Isaac Sim Replicator Composer is employed to capture the instance segmentation mask of the entire scene from a viewpoint (Figure 6(b)). Subsequently, each object’s visible mask is cropped from the instance segmentation mask of the scene. To obtain the amodal mask of each object on the simulated tabletop scene, we have developed the subsequent steps.
Initially, all objects’ visibility is disabled. For each object $o$ within the scene, its visibility is enabled, and the instance segmentation function is utilized to capture its amodal mask (Figure 6(c)). Following this, we compute the object’s occlusion mask and occlusion rate, as presented in (Figure 6(d)). The occlusion mask of an object $o$ can be acquired by subtracting its visible mask from its amodal mask.
The occlusion rate of the object $o$ can be computed by dividing the number of pixels in the occlusion mask by the number of pixels in the amodal mask. If the occlusion rate of the object $o$ is equal to 1, it implies that object $o$ is completely obscured from the viewpoint, thus we do not save the object $o$’s annotation for this view. The visibility of object $o$ is then disabled to capture the masks of the next object. Following the preservation of all objects’ masks, we use Algorithm 1 to generate the Occlusion Order Adjacency Matrix (OOAM) for this viewpoint (Figure 6(e)). For a scene with $M$ objects, the OOAM contains $M\times M$ elements, where the element ($i$, $j$) is a binary value in the matrix which indicates whether object $i$ occludes object $j$. Given the OOAM, we can easily construct the Occlusion Order Directed Graph (OODG) to visualize the occlusion order in the viewpoint (Figure 6(e)). We provide a detailed explanation of the OODG in our supplementary materials. After that, the visibility of all objects is enabled to prepare for the capturing of annotations from the next viewpoint of the scene.
@@ 4 Dataset Details
To demonstrate the capabilities of SynTable, we generated a sample synthetic dataset, SynTable-Sim, using our pipeline, for training and evaluating UOAIS models. Note that users can also generate other custom datasets that meet the specific requirements of their application using SynTable.
Table 1: A comparison of publicly available unseen object instance segmentation datasets for cluttered tabletop scenes. ##\## indicates the number of items. VI: Visible Instances. OI: Occluded Instances. Avg. OR %: Avg. Occlusion Rate %, i.e., the fraction of occluded pixels to amodal pixels across all object instances in the dataset. AM: Availability of amodal masks. OM: Availability of occlusion masks. Order: Availability of occlusion order relation information between objects. R/S: Real or Synthetic. - indicates that the data was not available in the literature. $*{}^{*}$ indicates that the values were not provided in the original literature, but we were able to compute the values.
Dataset
#Images
#Objects
#Scenes
#VI
#OI
Avg. OR (%)
AM
OM
Order
R/S
OCID [24]
2,390
89
96
R
OSD [23]
111
111
R
OSD-Amodal [2]
111
111
$474^{*}$
R
UOAIS-Sim
25,000 375 500 $356,885^{*}$ $127,129^{*}$ $11.16^{*}$ S

### v3

@@ Abstract
In this work, we present SynTable, a unified and flexible Python-based dataset generator built using NVIDIA’s Isaac Sim Replicator Composer for generating high-quality synthetic datasets for unseen object amodal instance segmentation of cluttered tabletop scenes. Our dataset generation tool can render complex 3D scenes containing object meshes, materials, textures, lighting, and backgrounds. Metadata, such as modal and amodal instance segmentation masks, object amodal RGBA instances, occlusion masks, depth maps, bounding boxes, and material properties can be automatically generated to annotate the scene according to the users’ requirements. Our tool eliminates the need for manual labeling in the dataset generation process while ensuring the quality and accuracy of the dataset. In this work, we discuss our design goals, framework architecture, and the performance of our tool. We demonstrate the use of a sample dataset generated using SynTable for training a state-of-the-art model, UOAIS-Net. Our state-of-the-art results show significantly improved performance in Sim-to-Real transfer when evaluated on the OSD-Amodal dataset. We offer this tool as an open-source, easy-to-use, photorealistic dataset generator for advancing research in deep learning and synthetic data generation. The links to our source code, demonstration video, and sample dataset can be found in the supplementary materials.
@@ 1 Introduction
Figure 1: (a) RGB outputs of photorealistic cluttered tabletop scenes generated by SynTable pipeline. (b) Visualization of RGB Images, Depth Images, Object Amodal Masks, Object Visible Masks, Object Occlusion Masks, and Object Visible Bounding Boxes.
Amodal completion is a perceptual ability that enables the perception of whole objects, even when they are partially occluded [16, 1]. It encompasses three key tasks: amodal shape completion, amodal appearance completion and occlusion order. Amodal shape completion involves predicting the complete structure of an object beyond its visible portion, typically represented as a binary segmentation mask that includes both visible and occluded regions. Amodal appearance completion refers to the process of inferring the likely apperance of the hidden regions of an object based on its visible parts (RGB values of hidden pixels). Occlusion Order considers the occlusion relationship between objects, distinguishing between occluders (objects that obscure others) and occludees (objects being occluded), which can involve no occlusion or bi-directional occlusion. Humans are capable of “filling in” the occluded appearance of invisible objects, owing to their vast experience in perceiving countless objects in various contexts and scenes. This ability to infer an object’s complete structure from its partial appearance is critical for systems requiring holistic scene understanding, such as augmented or virtual reality, and robotics and automation. In modern vision systems, accurately comprehending occluded objects in cluttered environments is essential for tasks ranging from object interaction to environment reconstruction.
There are three key challenges in amodal instance segmentation: Firstly, the lack of large-scale, high-quality datasets for unseen object amodal instance segmentation (UOAIS) limits the performance of vision systems in real-world applications [3]. While datasets exist for object detection and segmentation [6, 14, 17, 30, 10], only a few address UOAIS [2]. This is largely due to the difficulty of manually annotating amodal data, as human annotators must estimate occluded regions, leading to inherent subjectivity and inconsistencies in ground-truth annotations [2, 22, 32].
Secondly, synthetic datasets often suffer from visual domain mismatch due to non-photorealistic rendering or insufficient domain randomization [29], resulting in poor Sim-to-Real transfer. Existing tools prioritize rendering speed over photorealism, limiting their utility for training robust vision models, which results in a poor Sim-to-Real transfer that will inevitably reduce the performance of algorithms in real-world applications.
Thirdly, the lack of automated tools for generating amodal annotations and evaluating occlusion relationships hinders progress in this domain. Existing evaluation metrics focus on visible object regions but do not assess a model’s ability to infer occlusion order — a critical capability for systems operating in cluttered scenes. For example, understanding occlusion hierarchies enables sequential task planning and reduces errors caused by overlapping objects. However, manual annotation of such relationships is prohibitively time-consuming, necessitating simulation tools as a more cost-effective and accurate solution.
@@
In this work, we address these challenges by developing SynTable, a unified Python-based tool for generating customizable, photorealistic datasets for UOAIS in cluttered scenes. While our experiments focus on tabletop environments (common in interaction tasks), our framework generalizes to diverse settings. SynTable integrates rendering and annotation into a single pipeline, allowing users to control scene complexity, object variety, and annotation types. Built on NVIDIA’s Isaac Sim Replicator Composer, it leverages high-fidelity ray tracing and domain randomization to bridge the Sim-to-Real gap.
@@
    We develop a pipeline to automatically render photorealistic cluttered tabletop scenes and generate ground truth amodal instance segmentation masks, eliminating manual labeling in dataset generation. Our designed dataset generation tool creates photorealistic and accurately-labeled custom datasets for UOAIS (refer to Figure 1(a)).
@@
    Our tool provides a rich set of annotations related to amodal instance segmentation (refer to Figure 1(b)): modal (visible) and amodal instance segmentation masks, RGBA object instances, occlusion masks, occlusion rates, and occlusion order adjacency matrix. Users can easily select which annotations to include in their dataset based on the requirements of their application.
@@
    We generated an open-sourced large-scale sample synthetic dataset using our tool consisting of amodal instance segmentation labels for users to train and evaluate amodal segmentation models on 1075 novel objects, designed to benchmark amodal segmentation in occlusion-rich scenarios.
@@ 2 Related Works
2.1 Amodal Instance Segmentation in Vision Systems
Recent advances in amodal instance segmentation aim to enhance object detection and tracking in complex scenes. However, challenges such as limited training data and Sim-to-Real gaps persist, particularly in cluttered environments where occlusion reasoning is critical.
Lack of Large-scale High-quality Training Data. While datasets like [9, 7, 31] have advanced amodal segmentation for indoor scenes, few address occlusion-rich scenarios in everyday interaction tasks. Existing efforts often focus on narrow domains: for example, [12] introduced a benchmark for multi-object interaction in industrial settings, but its limited scene and object diversity restrict broader applicability. Similarly, the Object Segmentation Database (OSD) [24] and Object Cluttered Indoor Dataset (OCID) [25] pioneered tools for segmentation in cluttered scenes but lack amodal annotations. Recent work by Back et al. [2] manually added amodal masks to OSD, yet this approach remains labor-intensive and prone to human error.
Sim-to-Real Problem. Synthetic datasets like the Tabletop Object Dataset (TOD) [29] and UOAIS-Sim [2] struggle with photorealism and domain randomization, leading to significant Sim-to-Real gaps. For instance, TOD’s non-photorealistic rendering limits its utility for training models deployed in real-world applications such as augmented reality or autonomous navigation.
@@ 2.2 Tools for Generating Synthetic Datasets
With the rapid development of deep learning, the demand of researchers for synthetic datasets has increased in recent years, leading to the increased development of various tools for generating these datasets [28]. For robotics and computer vision applications, PyBullet and MuJoCo [27] are commonly used physical simulators to generate synthetic data. Xie et al. [29] pre-trained an RGB-D unseen object instance segmentation model using PyBullet. Tobin et al. [26] used MuJoCo to generate synthetic images with domain randomization, which can bridge the Sim-to-Real gap by realistically randomizing 3D content. Simulation tools such as PyBullet and MuJoCo typically come with renderers that are accessible and flexible, but they lack physically based light transport simulation, photorealism, material definitions, and camera effects.
@@ 3 Method
Our dataset generation pipeline is illustrated in Figure 2. Parameters and configurations of the scenes to be rendered are defined in a parameter file. Objects, materials, and light sources used in our pipeline are referred to as assets. The scene is prepared by rendering a tabletop scene with floating objects in Isaac Sim. A physical simulation is run to drop the rendered objects onto the table. For every view within a scene, camera viewpoints and lighting conditions are re-sampled. Subsequently, the annotations are captured to create the dataset. We provide additional details about each step of our data generation pipeline in Section 8 of our supplementary materials.
@@ 3.1 Preparing Each Scene
To prepare each scene, a table is randomly sampled and rendered in the center of a room, as shown in Figure 4. The texture and materials of the table, ceiling, wall, and floor are randomized for domain randomization while objects are added with randomized coordinates and orientations. We randomly sample (with replacement) $N_{lower}$ to $N_{upper}$ number of objects for each scene. Objects are initialized with real-life dimensions, mass, collision properties, randomized rotations and coordinates, ensuring diverse object arrangements across scenes. Additional details about our scene preparation method can be found in Section 8.1.
@@ 3.2 Physical Simulation of Each Scene
Rendered objects are dropped onto the table through a physics simulation to ensure the random placement of objects in the scene. Objects that rebound off the tabletop surface and land beyond the spatial coordinate region of the tabletop surface are removed, excluding extraneous objects from annotations. We provide more details about our physical simulation in Section 8.2.
@@ 3.3 Sampling of Camera Viewpoints
To capture annotations for each scene from multiple viewpoints, we enhance the approach of Gilles et al. [18] (which only uses fixed viewpoints) by capturing the $V$ number of viewpoints at random positions within custom radii of two concentric hemispheres of custom radii. The calculation of the Cartesian coordinates of each viewpoint can be found in 8.3 of our supplementary materials. Each viewpoint is oriented such that the camera looks directly at the center of the tabletop surface.
@@ 3.4 Sampling of Lighting Conditions
To simulate various indoor lighting conditions for each viewpoint, we resample $L$ spherical light sources using a method similar to Section 8.3. Please refer to Section 8.4 in our supplementary materials for more details. In contrast to Back et al.’s [2] approach of using point light sources, we use spherical light sources emitting light in all directions to mimic light bulbs. Furthermore, we uniformly sample the temperatures and intensity of the light sources. Users can customize the number of spherical light sources, as well as their intensities and temperatures.
Figure 3: The process of capturing annotations for a scene. For each viewpoint, (a) RGB and depth with all objects (b) object visible masks & bounding box, (c) object amodal masks (including object amodal RGBA instances), (d) object occlusion masks and occlusion rate, (e) occlusion order adjacency matrix are captured.
@@ 3.5 Capturing of Ground Truth Annotations
The process of capturing the annotations for a scene is illustrated in Figure 3. In each view, the RGB and depth images of the tabletop scene will be captured (Figure 3(a)). The built-in segmentation function in Isaac Sim Replicator Composer is used to capture the scene’s instance segmentation mask from a viewpoint (Figure 3(b)). Subsequently, the visible mask of each object is cropped from the scene’s segmentation mask.
For object amodal mask generation, we have developed the following steps. Initially, all objects’ visibility are disabled. For each object $o$ in the scene, its visibility is enabled and the instance segmentation function is utilized to capture its amodal mask and the amodal RGBA instance (Figure 3(c)). We compute the object’s occlusion mask and occlusion rate, as presented in (Figure 3(d)). After capturing all object masks, we use Algorithm 1 to generate the Occlusion Order Adjacency Matrix (OOAM) for this viewpoint (Figure 3(e)). For a scene with $M$ objects, the OOAM contains $M\times M$ elements, where the element ($i$, $j$) is a binary value in the matrix that indicates whether the object $i$ occludes the object $j$. Given the OOAM, we can easily construct the Occlusion Order Directed Graph (OODG) to visualize the occlusion order in the viewpoint (Figure 3(e)). We provide a detailed explanation of the OODG in Section 11 of our supplementary materials. After that, the visibility of all objects is enabled to prepare for the capturing of annotations from the next viewpoint of the scene.
@@ 4 Dataset Details
To demonstrate the capabilities of SynTable, we generated a sample synthetic dataset of cluttered tabletop scenes, SynTable-Sim, using our pipeline, to train and evaluate UOAIS models. Note that users can also generate other custom datasets that meet the specific requirements of their application using the SynTable pipeline.
Table 1: A comparison of publicly available unseen object instance segmentation datasets for cluttered tabletop scenes. ##\## indicates the number of items. VI: Visible Instances. OI: Occluded Instances. Avg. OR %: Avg. Occlusion Rate %, i.e., the fraction of occluded pixels to amodal pixels across all object instances in the dataset. AM: Availability of amodal masks. OM: Availability of occlusion masks. Order: Availability of occlusion order relation information between objects. R/S: Real or Synthetic. - indicates that the data was not available in the literature. indicates that the values were not provided in the original literature, but we were able to compute the values.
Dataset #Images #Objects #Scenes #VI #OI Avg. OR (%) AM OM Order R/S
OCID [25] 2,390 89 96 19,097 - - R
OSD [24] 111 - 111 474 - - R
OSD-Amodal [2] 111 - 111 $474^{*}$ 237 24.11 R
UOAIS-Sim 25,000 375 500 $356,885^{*}$ $127,129^{*}$ $11.16^{*}$ S